# Data-Driven Logistic Regression Ensembles With Applications in Genomics

### v5

@@ 7 The Number of Models
The analysis of genomics data via high-throughput technologies has generated the need for classification algorithms that can handle high-dimensional data containing correlated predictors (genes) within different pathways or networks, see Yousefi et al. (2011) and Zhang and Coombes (2012) for example. In light of this, to investigate the accuracy-diversity trade-off of split logistic regression, we use the high-dimensional block correlation setting of Scenario 3 in the previous section with configuration parameters $(n,p)=(50,\,$1,000), $(\rho_{1},\rho_{2})=(0.2,0.5)$, $\zeta\in\{0.1,0.2,0.4\}$ and $\mathbb{P}(Y=1)=0.2$.
@@ 7.1 Results
Table 6 shows the evolution of the ensemble misclassification rate (MR), the average misclassification rate of the individual models ($\overline{\text{MR}}$) and the entropy diversity measure (EM) averaged over the test sets, as a function of the number of models for split logistic regression. It also contains the overlap (OVP) between the individual models in the ensemble, defined as
@@
It can be seen that in all three settings $\overline{\text{MR}}$ and EM increase with the number of models, while the ensemble MR decreases. Hence, as the number of models increases, the accuracy of the individual models has less impact on the ensemble MR compared to their level of diversity. Split logistic regression manages to achieve a proper balance for this trade-off, resulting in a low MR for the ensemble.
Table 6: MR, ${\overline{\text{MR}}}$, EM and OVP as a function of the number of models under Scenario 3 averaged over the test sets.
@@
=2pt $\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$ $\mathbf{G}$ MR $\mathbf{\overline{MR}}$ EM OVP MR $\mathbf{\overline{MR}}$ EM OVP MR $\mathbf{\overline{MR}}$ EM OVP 1 0.14 -- -- -- 0.12 -- -- -- 0.12 -- -- -- 2 0.13 0.14 0.05 0.76 0.12 0.12 0.06 0.70 0.11 0.12 0.04 0.77 5 0.13 0.14 0.10 0.54 0.11 0.12 0.13 0.36 0.10 0.12 0.11 0.37 7 0.13 0.14 0.11 0.48 0.11 0.12 0.13 0.34 0.10 0.12 0.13 0.24 10 0.12 0.15 0.14 0.31 0.11 0.13 0.14 0.21 0.10 0.12 0.13 0.19 15 0.12 0.16 0.20 0.19 0.10 0.14 0.19 0.09 0.10 0.13 0.17 0.09 20 0.12 0.17 0.21 0.11 0.10 0.14 0.20 0.05 0.10 0.13 0.18 0.05 25 0.12 0.17 0.24 0.09 0.10 0.14 0.22 0.05 0.09 0.14 0.20 0.05
@@ 7.2 Computational Cost
Table 6 indicates that a larger number of models results in an ensemble with lower misclassification rate. However, MR stabilizes quickly, so there is a diminishing returns type of behavior in terms of prediction accuracy versus computational cost. Indeed, we also ran split logistic regression using $G=50$ models, but in all cases there is hardly any improvement in MR compared to the ensemble with $G=25$ models shown in Table 6. In fact, with $G=25$ models split logistic regression already achieves nearly full diversity (OV $\approx 1/G$), so little gain is expected by increasing the number of models further while computation time does grow.
Table 7.2 shows the average computation time (in CPU seconds) across all sparsity levels of scenario 3 as a function of the number of models. As $G$ increases there is a price to pay in average individual ${\overline{\text{MR}}}$ which may not be compensated by a lower ensemble MR. The computation time seems to depend linearly on the number of models and is approximately given by $4.88+1.27\times G$ for $G\geq 2$. In real data applications such as the gene expression applications in the previous section it is generally a good strategy to use a data-driven choice of the number of models in the ensemble by increasing $G$ until the CV performance has stabilized.
Table 7: Computation time of R function call for split logistic regression in CPU seconds for varying number of models. CPU seconds are on a 2.7 GHz Intel Xeon processor in a machine running Linux 7.8 with 125 GB of RAM.
@@ 8 Discussion and Future Directions
Due to the diversity penalty, split logistic regression makes use of different groups of variables in the individual models to build an ensemble. Allowing interactions among predictors can be beneficial to further improve the prediction performance of classifiers. Since split logistic regression can have much higher recall than single-model methods such as lasso and the elastic net, our methodology can also be useful to detect interaction effects that would be missed by such single-model methods. This is important for example for genomics data applications where it is known that gene interaction effects are common. The diversity penalty can also be combined more generally with different sparsity penalties such as the group lasso (Meier et al., 2008) for categorical variables or the fused lasso (Tibshirani et al., 2005) for data exhibiting spatial or temporal structures.
@@
Ensemble methods are very popular to analyze small sample data with a large number of predictor variables, and our method provides a framework to build an optimal classification ensemble model. Similarly to logistic regression, the general split modeling framework could be applied to multi-class classification problems to obtain a powerful ensemble classifier. The split modeling framework could also be extended to generalized linear models in general.
@@ Acknowledgments
The supplementary material contains the consistency proof, details of our algorithm, the analysis with the alternative diversity measures and the full results of our simulation and medical genomics data experiments. The R package SplitGLM along with its reference manual is publicly available on CRAN at https://CRAN.R-project.org/package=SplitGLM. The data and scripts to replicate the numerical experiments are available at https://doi.org/10.5281/zenodo.10250506.
@@ 8.11 Results
In this section, the full results of the simulation study in Section 5 of the article are reported. For the three scenarios described in the article and the considered values of the correlation parameters ($\rho$, $\rho_{1}$ and $\rho_{2}$), sample size ($n$), sparsity level ($\zeta$), and probability of positive events ($\pi_{1}$), the tables report the misclassification rate (MR), sensitivity (SE), specificity (SP), test-sample loss (TL), recall (RC) and precision (PR) for the eleven methods listed in Section 5.2 of the article.
@@
    • Tables 2-7 contain the results for the MR, SE and SP.
    • Tables 8-13 contain the results for the TL, RC and PR.
@@
    • Tables 14-19 contain the results for the MR, SE and SP.
    • Tables 20-25 contain the results for the TL, RC and PR.
@@
    • Tables 26-31 contain the results for the MR, SE and SP.
    • Tables 32-37 contain the results for the TL, RC and PR.

### v6

@@ 7 The Number of Models
The analysis of genomics data via high-throughput technologies has generated the need for classification algorithms that can handle high-dimensional data containing correlated predictors (genes) within different pathways or networks, see Yousefi et al. (2011) and Zhang and Coombes (2012) for example. In light of this, to investigate the accuracy-diversity trade-off of split logistic regression, we use the high-dimensional block correlation setting of Scenario 3 in Section 5 with configuration parameters $(n,p)=(50,\,$1,000), $(\rho_{1},\rho_{2})=(0.2,0.5)$, $\zeta\in\{0.1,0.2,0.4\}$ and $\mathbb{P}(Y=1)=0.2$.
@@ 7.1 Results
Table 10 shows the evolution of the ensemble prediction accuracy (ACC), the average prediction accuracy of the individual models ($\overline{\text{ACC}}$) and the entropy diversity measure (EM) averaged over the test sets, as a function of the number of models for split logistic regression. It also contains the overlap (OVP) between the individual models in the ensemble, defined as
@@
It can be seen that in all three settings $\overline{\text{ACC}}$ decreases and EM increases with the number of models, while the ensemble ACC increases. Hence, as the number of models increases, the accuracy of the individual models has less impact on the ensemble ACC compared to their level of diversity. Split logistic regression manages to achieve a proper balance for this trade-off, resulting in a high ACC for the ensemble.
@@
[…]
=2pt $\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$ $\mathbf{G}$ ACC $\mathbf{\overline{ACC}}$ EM OVP ACC $\mathbf{\overline{ACC}}$ EM OVP ACC $\mathbf{\overline{ACC}}$ EM OVP 1 0.86 -- -- -- 0.87 -- -- -- 0.87 -- -- -- 2 0.87 0.86 0.04 0.81 0.89 0.88 0.05 0.76 0.88 0.88 0.05 0.73 5 0.87 0.86 0.10 0.53 0.89 0.88 0.12 0.44 0.89 0.88 0.11 0.42 7 0.87 0.86 0.10 0.47 0.90 0.88 0.14 0.32 0.89 0.88 0.11 0.39 10 0.87 0.85 0.13 0.29 0.90 0.87 0.16 0.17 0.90 0.87 0.14 0.19 15 0.87 0.84 0.18 0.20 0.90 0.86 0.20 0.12 0.90 0.87 0.18 0.07 20 0.87 0.84 0.19 0.12 0.90 0.86 0.21 0.08 0.90 0.86 0.19 0.05 25 0.87 0.83 0.20 0.13 0.90 0.85 0.24 0.04 0.90 0.86 0.21 0.04
@@ 7.2 Computational Cost
Table 10 indicates that a larger number of models results in an ensemble with higher prediction accuracy. However, ACC stabilizes quickly, so there is a diminishing returns type of behavior in terms of prediction accuracy versus computational cost. Indeed, we also ran split logistic regression using $G=50$ models, but in all cases there is hardly any improvement in ACC compared to the ensemble with $G=25$ models shown in Table 10. In fact, with $G=25$ models split logistic regression already achieves nearly full diversity (OV $\approx 1/G$), so little gain is expected by increasing the number of models further while computation time does grow.
Table 7.2 shows the average computation time (in CPU seconds) across all sparsity levels of scenario 3 as a function of the number of models. As $G$ increases there is a price to pay in average individual ${\overline{\text{ACC}}}$ which may not be compensated by a higher ensemble ACC. The computation time seems to depend linearly on the number of models and is approximately given by $0.27+0.15\times G$ for $G\geq 2$. In real data applications such as the gene expression applications in the previous section it is generally a good strategy to use a data-driven choice of the number of models in the ensemble by increasing $G$ until the CV performance has stabilized.
Table 11: Computation time of R function call for split logistic regression in CPU seconds for varying number of models, using multithreading (5 threads). CPU seconds are on a 2.1 GHz Intel Xeon Platinum 8468 processor in a machine running CentOS Linux 7.9 with 32 GB of RAM.
@@ 8 Discussion and Future Directions
While this manuscript focuses on microarray gene expression data, which has been a canonical example of the $p\gg n$ scenario in statistical genomics for decades, the proposed split logistic regression approach is equally applicable to other types of high-dimensional omics data. Modern technologies like RNA-seq, single-cell sequencing, proteomics, and metabolomics all generate data with similar dimensional characteristics. For instance, single-cell RNA sequencing data can be transformed into pseudobulk profiles representing cell populations, methylation data can be summarized at the gene or region level, and proteomics datasets frequently contain thousands of measured proteins across limited samples. The common thread across these data types is that they all present the fundamental statistical challenge of extracting meaningful signals from thousands of features measured on relatively few subjects—precisely the scenario where our proposed method demonstrates advantages over traditional approaches.
Due to the diversity penalty, split logistic regression makes use of different groups of variables in the individual models to build an ensemble. Allowing interactions among predictors can be beneficial to further improve the prediction performance of classifiers. Since split logistic regression can have much higher recall than single-model methods such as Lasso and the elastic net, the proposed methodology can also be useful to detect interaction effects that would be missed by such single-model methods. This is important for example for genomics data applications where it is known that gene interaction effects are common. The diversity penalty can also be combined more generally with different sparsity penalties such as the group Lasso (Meier et al., 2008) for categorical variables or the fused Lasso (Tibshirani et al., 2005) for data exhibiting spatial or temporal structures.
@@
Ensemble methods are very popular to analyze small sample data with a large number of predictor variables, and the proposed method provides a framework to build an optimal classification ensemble model. Similarly to logistic regression, the general split modeling framework could be applied to multi-class classification problems to obtain a powerful ensemble classifier. The split modeling framework could also be extended to generalized linear models in general.
@@ Acknowledgments
The supplementary material contains the consistency proof, details of our algorithm, the analysis with the alternative diversity measures and the full results of our simulation and medical genomics data experiments. The R/C++ package SplitGLM along with its reference manual is publicly available on CRAN at https://CRAN.R-project.org/package=SplitGLM. The data and scripts to replicate the numerical experiments are available at https://doi.org/10.5281/zenodo.15588653.
@@ 8.11 Results
In this section, the full results of the simulation study in Section 5 of the article are reported. For the five scenarios described in the article and the considered values of the correlation parameters ($\rho$, $\rho_{1}$ and $\rho_{2}$), sample size ($n$), sparsity level ($\zeta$), and probability of positive events ($\pi_{1}$), the tables report the prediction accuracy (ACC), sensitivity (SNS), specificity (SPC), area under the ROC Curve (AUC), test-sample loss (TSL), recall (RCL) and precision (PRC) for the fourteen methods listed in Section 3 of the article.
@@
    • Tables 2-7 contain the results for the ACC, SNS, SPC, AUC and TSL.
    • Tables 8-13 contain the results for the RCL and PR.
@@
    • Tables 14-19 contain the results for the ACC, SNS, SPC, AUC and TSL.
    • Tables 20-25 contain the results for the RCL and PR.
@@
    • Tables 26-31 contain the results for the ACC, SNS, SPC, AUC and TSL.
    • Tables 32-37 contain the results for the RCL and PR.
  • Scenario 4:
    • Tables 38-43 contain the results for the ACC, SNS, SPC, AUC and TSL.
    • Tables 44-49 contain the results for the RCL and PR.
  • Scenario 5:
    • Tables 50-55 contain the results for the ACC, SNS, SPC, AUC and TSL.
    • Tables 56-61 contain the results for the RCL and PR.